# TGB 2.0: A Benchmark for Learning on Temporal Knowledge Graphs and Heterogeneous Graphs

### v1

@@ 1 Introduction
Learning from graph-structured data has become ubiquitous in many applications such as recommendation systems [33, 66], knowledge base completion [52, 43] and molecular learning [55, 3]. Relational data often evolves over time and can contain multiple types of relations. These complex interactions and temporal dependencies can be captured by multi-relational temporal graphs. In recent years, various approaches have emerged to predict future links in such graphs, notably for prediction on Temporal Knowledge Graphs (TKGs) [39, 44] and Temporal Heterogeneous Graphs (THGs) [36, 28]. These approaches capture the rich information from multi-relational data, developing distinct lines of research from that of single-relational temporal graphs [58, 47]. However, benchmarking on multi-relational temporal graphs faces two main challenges: inconsistent evaluation and limited dataset size.
@@
Figure 1: Existing benchmark datasets (yellow) vs. novel datasets in TGB 2.0 for TKG (a) marked in orange and THG (b) marked in green. Circle sizes correspond to the number of timestamps. TGB 2.0 datasets are signficantly larger than existing datasets in number of nodes, edges and timestamps.
@@
Limited Dataset Size. Existing evaluations are conducted on predominantly small-scale datasets. For example, commonly used TKG and THG datasets consists of less than two million edges and one million nodes [36, 28, 39, 37]. However, real world networks typically contains tens of millions of nodes and edges thus existing datasets rarely reflect the true scale of datasets in practice. In addition, significant efforts were made to design scalable graph learning methods for real applications which requires the availability of large scale datasets [25, 24, 26]. These challenges hinder meaningful comparisons between methods and the accurate assessment of progress, hamper advancements in the field. Therefore, there is an urgent need for a public and standardized benchmark to facilitate proper and fair comparison between methods, accelerating research for multi-relational temporal graphs.
To address the aforementioned challenges, we present TGB 2.0, a novel benchmark designed for future link prediction on multi-relational temporal graphs. Building upon the foundations of the Temporal Graph Benchmark (TGB) [27] where only single-relation temporal graphs are included, TGB 2.0 introduces multi-relational temporal graph datasets. TGB 2.0 adds four novel TKG datasets and four novel THG datasets of varying scale, spanning five domains. Figure 1 shows the difference in scale of the novel datasets in TGB 2.0 when compared to existing ones. Figure 1(a) shows that TGB 2.0 TKG datasets (marked in orange) are orders of magnitude larger than existing ones in terms of number of nodes, edges and timestamps. Figure 1(b) shows that TGB 2.0 THG datasets (marked in green) are significantly larger than existing datasets. With thgl-myket dataset quintupling the number of edges and timestamps while thgl-github has the most number of nodes to date. Additionally, TGB 2.0 provides an automated evaluation pipeline for reproducible and realistic evaluation on multi-relational temporal graphs. In TGB 2.0, the dynamic link property prediction task is treated as a ranking problem where multiple negative edges are ranked against the positive edge. For large datasets, we sample negative edges based on the edge type of the query, thus closely approximates the complete evaluation where all negative edges are used. Overall, TGB 2.0 presents a benchmark for realistic, challenging, and reproducible evaluation on multi-relational temporal graphs while providing an automated pipeline for dataset downloading, processing, evaluation as well as a public leaderboard. TGB 2.0 has the following main contributions:
@@
Reproducibility. TGB 2.0 code and datasets are publicly available (see Appendix C for download links) and theTGB 2.0 website provides detailed documentation.
@@ 2 Preliminaries
Time Representations. There are two approaches for representing time in temporal graphs: (a) discrete, where graphs are modeled as snapshots $G_{t}$ containing all edges appearing at time $t$, and (b) graphs that can be conceptualized as a series of edges arriving at continuous timestamps. In practice, TKGs are often represented as snapshots and TKG methods are tailored for discrete time representations [39, 37, 44, 63]. This choice is driven by the discrete nature of the data sources and the suitability of snapshot-based representations for downstream tasks. Therefore, we represent TKGs as snapshots in TGB 2.0. In comparison, THG data sources are often continuous in nature (recorded in second-wise interactions), while both discrete and continuous approaches for THGs are developed [36, 8, 68]. It is often argued that continuous representations preserves more information and can be converted to discrete while the reverse is not true [31]. Therefore, the THG datasets are represented in the continuous format.
@@ 3 Related Work
TKG Methods. Most TKG forecasting methods utilize a discrete time representation, except for [22]. Some methods integrates the message-passing paradigm from static graphs [60, 50] with sequential techniques [30, 39, 20, 21, 37, 42]. Other approaches combine Reinforcement Learning with temporal reasoning for future link prediction [38, 63]. Rule-based methods [44, 32, 49, 45, 41] employ strategies to learn temporal logic rules while others [72, 67, 71] combines a blend of different methodologies. More details are in Appendix G.1. For TKG forecasting, common benchmark datasets include YAGO [48], WIKI [34, 29], GDELT [35] and the Integrated Crisis Early Warning System (ICEWS) dataset [6]. However, these datasets are orders of magnitude smaller than our TKG datasets in number of nodes, edges and timestamps. In tkgl-icews, we include the full ICEWS dataset [6] spanning 28 years when comparing to prior versions containing only one or a few years [12, 29, 9]. Similarly, our tkgl-wikidata dataset is orders of magnitude larger than the existing WIKI dataset [13, 40] in size of nodes, edges and timestamps.
THG Methods. THG methods can be categorized based on their time representation: discrete-time methods and continuous-time methods. Examples of continous time methods include HTGN-BTW [70] and STHN [36]. HTGN-BTW [70], enabling TGN [58] to accommodate heterogeneous node and edge types. STHN [36] utilizes a link encoder and patching techniques to incorporate edge type and time information respectively. Discrete-time methods includes random walk based methods [5, 69] and message-passing based methods[68, 8]. However, it is difficult to adapt discrete-time methods for continuous-time datasets. More details are provided in Appendix G.2. Common THG datasets such as MathOverflow [53], Netflix [4] and Movielens [23] are small and only contain a few million edges [36]. Large datasets such as Dataset A and B from the WSDM 2022 Challenge and the TRACE and THEIA datasets [56] are only evaluated with one negative sample per positive edge, which is shown to lead to over-optimistic and insufficient evaluation [26, 54]. Here, we introduce the large thgl-myket dataset with 53 million edges and 14 million timestamps.
@@ 4 Datasets
TGB 2.0 introduces eight novel datasets from five distinct domains consisting of four TKGs and four THGs. We split all datasets chronologically into training, validation, and test sets, respectively containing $70\%$, $15\%$, and $15\%$ of all edges in line with existing studies [26, 58, 47] and ensure that edges of a timestamp can only exist in either train or validation or test set [^1] We detail the exact number of timestamps and edges for each subset in Appendix F.6. We present the dataset licenses and download links in Appendix D. The datasets will be permanently maintained via Digital Research Alliance of Canada (funded by the Government of Canada).
Dataset Details. Here we describe each TGB 2.0 dataset in detail. Temporal Knowledge Graph datasets start with the prefix tkgl- while Temporal Heterogeneous Graph datasets start with thgl-.
@@ 5 Experiments
1-vs-all. For datasets where there is a small number of nodes, it is possible to evaluate with all the possible destinations, thus achieving a comprehensive evaluation. In TGB 2.0, we use 1-vs-all strategy for tkgl-smallpedia, tkgl-polecat and tkgl-icews due to their smaller node size (see Table 1).
@@ 5.1 Temporal Knowledge Graph Experiments
Table 2: MRR results for Temporal Knowledge Graph link property prediction task. We report the average and standard deviation across 5 different runs. First place is bolded and second place is underlined.
@@
$\text{RecB}_{\text{train}}$ [13] 0.694 0.655 0.203 0.198 OOT OOT OOT OOT
$\text{RecB}_{\text{default}}$ [13] 0.640 0.570 0.170 0.167 0.264 0.206 OOT OOT
@@
Figure 4: MRR per relation for the 10 highest occuring relations for three TKG datasets for RE-GCN, CEN and Reccurency Baseline. The color indicates the Recurrency Degree value for relation type. The relations for each dataset are ordered by decreasing Recurrency Degrees.
@@
Per-relation Analysis. Figure 4 illustrates the performance per relation of selected methods across three datasets  [^2] The relation description can be found based on their relation ID in Figure 7 in Appendix.. For each dataset we show the ten most frequent relations, ordered by decreasing Recurrency Degree with the color reflecting the Recurrency Degree of each relation. Note that the y-axis scale varies across datasets. We observe distinct patterns in relation-specific performance across datasets: while results on tkgl-polecat exhibit consistent performance levels across relations, suggesting a relative homogeneity, results on the tkgl-smallpedia dataset shows significant variance in performance, indicating a higher degree of variation among relations. Interestingly, there is strong correlation between the Recurrency Degree and performance, most evident within the tkgl-smallpedia dataset.
@@ 5.2 Temporal Heterogenous Graph Experiments
Table 3: MRR results for Temporal Heterogeneous Graph Link Prediction task. We report the average and standard deviation across 5 different runs. First place is bolded and second place is underlined.
@@
For THG experiments, we include TGN [58] (with and without edge type information), STHN [36], RecB [13], and EdgeBank [54] for comparison. Table 3 reports the average performance and standard deviation across 5 runs for each method. We observe that scalability is a significant challenge for THG methods on large datasets such as thgl-forum and thgl-myket, more details can be found in Appendix F. Most methods either are out of memory or out of time for these datasets. STHN achieves the highest performance on thgl-software dataset showing method designed for THG can achieve significant performance gain. However, STHN is the least scalable, requiring 185 GB of memory for thgl-software to compute subgraphs and unable to scale to other datasets. The widely-used TGN model [58] for single-relation temporal graph learning is also adapted here, with a modification where it incorporates the edge type information as edge feature. We observe significant improvement when TGN utilizes edge type data, thus highlighting the potential to leverage the multi-relational information in THGs. Lastly, the EdgeBank heuristic achieves competitive performance with that of TGN while being scalable to large datasets. Therefore, it is important to evaluate against simple baselines to understand method performances.
@@ 6 Conclusion
In this work, we introduce TGB 2.0, a novel benchmark for reproducible, realistic, and robust evaluation on multi-relational temporal graphs that is building on the Temporal Graph Benchmark (TGB). We present four new TKG datasets and four new THG datasets which introduce multi-relation datasets in TGB. TGB 2.0 datasets are significantly larger than existing ones while being diverse in statistics and dataset domain. TGB 2.0 focuses on the dynamic link property prediction tasks and provides an automated machine learning pipeline for dataset downloading, processing, method evaluation, and a public leaderborad to track recent progress. From our experiments, we find that both TKG and THG methods struggle to tackle large scale datasets in TGB 2.0, often resulting in overly long runtime or exceeding the memory limit. Therefore, scalability is an important future direction. Another observation is that heuristic methods achieve competitive results on TKG and THG datasets. This highlights the importance of the inclusion of simple baselines and underlines the room for improvement for current methods.
@@ Acknowledgment
The authors extend their gratitude to Myket Corporation for generously providing the anonymous interaction data utilized in this research. We are thankful to Dr. Vahid Rahimian for agreeing to collaborate and for their support throughout the data-sharing process and to Ms. Zahra Eskandari for her diligent efforts in collecting and arranging the data. Additionally, we thank MohammadAmin Fazli for his continuous support and help in facilitating collaboration and data sharing. We thank Weihua Hu, Matthias Fey, Jure Leskovec and Michael Bronstein from the TGB team for their support and discussion. We thank the OGB team for sharing their website template for the TGB website. This research was enabled in part by compute resources, software and technical help provided by Mila. This research was supported by the Canadian Institute for Advanced Research (CIFAR AI chair program), Natural Sciences and Engineering Research Council of Canada (NSERC) Postgraduate Scholarship Doctoral (PGS D) Award and Fonds de recherche du Québec - Nature et Technologies (FRQNT) Doctoral Award.
Appendix A Limitations
This work exclusively considers the continuous-time setting for THG datasets. Depending on the application, either the continuous-time or discrete-time setting may be more appropriate. However, the continuous-time setting is often regarded as the more general framework. Nonetheless, many THG methods are designed for discrete settings. Thus, as future work, discretized versions of the datasets for comparative analysis between discrete methods could be added.
Additionally, the TGB 2.0 dataset collection currently includes datasets from only five distinct domains. Notably, domains such as biological networks and citation networks are not represented. To address this limitation, we plan to expand the dataset collection by incorporating additional datasets based on community feedback, thereby enhancing the diversity and comprehensiveness of the dataset repository.
Appendix B Broader Impact

### v2

@@ 1 Introduction
Learning from graph-structured data has become ubiquitous in many applications such as recommendation systems [35, 70], knowledge base completion [55, 46] and molecular learning [58, 3]. Relational data often evolves over time and can contain multiple types of relations. These complex interactions and temporal dependencies can be captured by multi-relational temporal graphs. In recent years, various approaches have emerged to predict future links in such graphs, notably for prediction on Temporal Knowledge Graphs (TKGs) [41, 47] and Temporal Heterogeneous Graphs (THGs) [38, 30]. These approaches capture the rich information from multi-relational data, forming distinct lines of research from those of single-relational temporal graphs [61, 50]. However, benchmarking on multi-relational temporal graphs faces two main challenges: inconsistent evaluation and limited dataset size.
@@
Figure 1: Existing benchmark datasets (yellow) vs. novel datasets in TGB 2.0 for TKG (a) marked in orange and THG (b) marked in green. Circle sizes correspond to the number of timestamps. TGB 2.0 datasets are significantly larger than existing datasets in number of nodes, edges and timestamps.
@@
Limited Dataset Size. Existing evaluations are conducted on predominantly small-scale datasets. For example, commonly used TKG and THG datasets consist of less than two million edges and one million nodes [38, 30, 41, 39]. However, real world networks typically contain tens of millions of nodes and edges thus existing datasets rarely reflect the true scale of datasets in practice. In addition, significant efforts were made to design scalable graph learning methods for real applications which requires the availability of large scale datasets [27, 26, 28]. These challenges hinder meaningful comparisons between methods and the accurate assessment of progress, and hamper advancements in the field. Therefore, there is an urgent need for a public and standardized benchmark to facilitate proper and fair comparison between methods, accelerating research for multi-relational temporal graphs.
To address the aforementioned challenges, we present TGB 2.0, a novel benchmark designed for future link prediction on multi-relational temporal graphs. Building upon the foundations of the Temporal Graph Benchmark (TGB) [29] where only single-relation temporal graphs are included, TGB 2.0 introduces multi-relational temporal graph datasets. TGB 2.0 adds four novel TKG datasets and four novel THG datasets of varying scale, spanning five domains. Figure 1 shows the difference in scale of the novel datasets in TGB 2.0 when compared to existing ones. Figure 1(a) shows that TGB 2.0 TKG datasets (marked in orange) are orders of magnitude larger than existing ones in terms of number of nodes, edges and timestamps. Figure 1(b) shows that TGB 2.0 THG datasets (marked in green) are significantly larger than existing datasets, with thgl-myket dataset quintupling the number of edges and timestamps, while thgl-github has the most number of nodes to date. Additionally, TGB 2.0 provides an automated evaluation pipeline for reproducible and realistic evaluation on multi-relational temporal graphs. In TGB 2.0, the dynamic link property prediction task is treated as a ranking problem where multiple negative edges are ranked against the positive edge. For large datasets, we sample negative edges based on the edge type of the query, and thus closely approximate the complete evaluation where all negative edges are used. Overall, TGB 2.0 presents a benchmark for realistic, challenging, and reproducible evaluation on multi-relational temporal graphs while providing an automated pipeline for dataset downloading, processing, evaluation as well as a public leaderboard. TGB 2.0 has the following main contributions:
@@
Reproducibility. TGB 2.0 code and datasets are publicly available (see Appendix B for download links) and theTGB 2.0 website provides detailed documentation.
@@ 2 Preliminaries
Time Representations. There are two approaches for representing time in temporal graphs: (a) discrete, where graphs are modeled as snapshots $G_{t}$ containing all edges appearing at time $t$, and (b) graphs that can be conceptualized as a series of edges arriving at continuous timestamps. In practice, TKGs are often represented as snapshots and TKG methods are tailored for discrete time representations [41, 39, 47, 66]. This choice is driven by the discrete nature of the data sources and the suitability of snapshot-based representations for downstream tasks. Therefore, we represent TKGs as snapshots in TGB 2.0. In comparison, THG data sources are often continuous in nature (recorded in second-wise interactions), while both discrete and continuous approaches for THGs are developed [38, 10, 72]. It is often argued that continuous representations preserve more information and can be converted to discrete while the reverse is not true [33]. Therefore, the THG datasets are represented in the continuous format.
@@ 3 Related Work
TKG Methods.
Most TKG forecasting methods utilize a discrete time representation, except for [24]. Some methods integrate the message-passing paradigm from static graphs [63, 53] with sequential techniques [32, 41, 22, 23, 39, 45]. Other approaches combine Reinforcement Learning with temporal reasoning for future link prediction [40, 66]. Rule-based methods [47, 34, 52, 48, 44] employ strategies to learn temporal logic rules while others [76, 71, 75] combines a blend of different methodologies. More details are in Appendix G.1. For more comprehensive discussions on methods and applications of TKG, we refer to the surveys [8][9][43], and [69]. For TKG forecasting, common benchmark datasets include YAGO [51], WIKI [36, 31], GDELT [37] and the Integrated Crisis Early Warning System (ICEWS) dataset [6]. However, these datasets are orders of magnitude smaller than our TKG datasets in number of nodes, edges and timestamps. In tkgl-icews, we include the full ICEWS dataset [6] spanning 28 years when comparing to prior versions containing only one or a few years [14, 31, 11]. Similarly, our tkgl-wikidata dataset is orders of magnitude larger than the existing WIKI dataset [15, 42] in size of nodes, edges and timestamps.
THG Methods. THG methods can be categorized based on their time representation: discrete-time methods and continuous-time methods. Examples of continous time methods include HTGN-BTW [74] and STHN [38]. HTGN-BTW [74] enables TGN [61] to accommodate heterogeneous node and edge types. STHN [38] utilizes a link encoder and patching techniques to incorporate edge type and time information respectively. Discrete-time methods include random walk based methods [5, 73] and message-passing based methods[72, 10]. However, it is difficult to adapt discrete-time methods for continuous-time datasets. More details are provided in Appendix G.2. Common THG datasets such as MathOverflow [56], Netflix [4] and Movielens [25] are small and only contain a few million edges [38]. Large datasets such as Dataset A and B from the WSDM 2022 Challenge and the TRACE and THEIA datasets [59] are only evaluated with one negative sample per positive edge, which is shown to lead to over-optimistic and insufficient evaluation [28, 57]. Here, we introduce the large thgl-myket dataset with 53 million edges and 14 million timestamps.
@@ 4 Datasets
TGB 2.0 introduces eight novel datasets from five distinct domains consisting of four TKGs and four THGs. We split all datasets chronologically into training, validation, and test sets, respectively containing $70\%$, $15\%$, and $15\%$ of all edges in line with existing studies [28, 61, 50] and ensure that edges of a timestamp can only exist in either train or validation or test set [^1] We detail the exact number of timestamps and edges for each subset in Appendix F.8. We present the dataset licenses and download links in Appendix C. The datasets will be permanently maintained via Digital Research Alliance of Canada (funded by the Government of Canada).
Dataset Details. Here we describe each TGB 2.0 dataset in detail. Temporal Knowledge Graph datasets start with the prefix tkgl- while Temporal Heterogeneous Graph datasets start with thgl-. More details on the dataset collection process are in Appendix D.
@@ 5 Experiments
1-vs-all. For datasets with a small number of nodes, it is possible to evaluate with all the possible destinations, thus achieving a comprehensive evaluation. In TGB 2.0, we use 1-vs-all strategy for tkgl-smallpedia, tkgl-polecat and tkgl-icews due to their smaller node size (see Table 1).
@@ 5.1 Temporal Knowledge Graph Experiments
Table 2: MRR results for Temporal Knowledge Graph Link Prediction task. We report the average and standard deviation across 5 runs. First place is bolded, second place underlined.
@@
$\text{RecB}_{\text{train}}$ [15] 0.639 0.605 0.203 0.198 0.270 0.211 OOT OOT
$\text{RecB}_{\text{default}}$ [15] 0.542 0.486 0.170 0.167 0.264 0.206 OOT OOT
@@
Figure 4: MRR per relation for the 10 highest occuring relations for three TKG datasets for RE-GCN, CEN and $\text{RecB}_{\text{train}}$. The color indicates the Recurrency Degree value for relation type. The relations for each dataset are ordered by decreasing Recurrency Degrees.
@@
Per-relation Analysis. Figure 4 illustrates the performance per relation of selected methods across three datasets  [^2] The relation description can be found based on their relation ID in Figure 7 in Appendix.. For each dataset, we show the ten most frequent relations, ordered by decreasing Recurrency Degree with the color reflecting the Recurrency Degree of each relation. Note that the y-axis scale varies across datasets. We observe distinct patterns in relation-specific performance across datasets: while results on tkgl-polecat exhibit consistent performance across relations, suggesting a relative homogeneity, results on the tkgl-smallpedia dataset shows significant variance, indicating a higher degree of variation among relations. Interestingly, there is strong correlation between the Recurrency Degree and performance, most evident within the tkgl-smallpedia dataset.
@@ 5.2 Temporal Heterogenous Graph Experiments
Table 3: MRR results for Temporal Heterogeneous Graph Link Prediction task. We report the average and standard deviation across 5 runs. First place is bolded, second place underlined.
@@
For THG experiments, we include TGN [61] (with and without edge type information), STHN [38], RecB [15], and EdgeBank [57]. Table 3 reports the average performance and standard deviation across 5 runs for each method. Scalability is a significant challenge for THG methods on large datasets such as thgl-forum and thgl-myket, more details can be found in Appendix F. Most methods either are out of memory or out of time for these datasets. STHN achieves the highest performance on thgl-software dataset showing methods designed for THG can achieve significant performance gain. However, STHN is the least scalable, requiring 185 GB of memory for thgl-software to compute subgraphs and unable to scale to other datasets. The widely-used TGN model [61] for single-relation temporal graph learning is also adapted here, with a modification to incorporate the edge type information as edge feature. We observe significant improvement when TGN utilizes edge type data, thus showing the potential to leverage the multi-relational information in THGs. Lastly, EdgeBank achieves competitive performance with that of TGN while being scalable to large datasets. Thus, it is important to evaluate against simple baselines to understand method performances.
@@ 6 Conclusion
In this work, we introduce TGB 2.0, a novel benchmark for reproducible, realistic, and robust evaluation on multi-relational temporal graphs that is building on the Temporal Graph Benchmark (TGB). We present four new TKG and four new THG datasets which introduce multi-relation datasets in TGB. TGB 2.0 datasets are significantly larger than existing ones while being diverse in statistics and dataset domains. TGB 2.0 focuses on the dynamic link property prediction task and provides an automated pipeline for dataset downloading, processing, method evaluation, and a public leaderboard to track progress. From our experiments, we find that both TKG and THG methods struggle to tackle large scale datasets in TGB 2.0, often resulting in overly long runtime or exceeding the memory limit. Therefore, scalability is an important future direction. Another observation is that heuristic methods achieve competitive results on TKG and THG datasets. This highlights the importance of the inclusion of simple baselines and underlines the room for improvement for current methods.
Limitations This work exclusively considers the continuous-time setting for THG datasets. Depending on the application, either the continuous-time or discrete-time setting may be more appropriate. However, the continuous-time setting is often regarded as the more general framework. However, many THG methods are designed for discrete settings. Thus, in future work, discretized versions of the data sets could be added for comparative analysis between discrete methods. Additionally, the TGB 2.0 dataset collection currently includes datasets from only five distinct domains. Notably, domains such as biological networks and citation networks are not represented. To address this limitation, we plan to expand the dataset collection by incorporating additional datasets based on community feedback, thereby enhancing the diversity and comprehensiveness of the dataset repository. As temporal graph data often records interactions between individuals or other sensitive information, data privacy is a potential concern. Sensitive information may be stored as node level or link level attributes thus proper anonymization of data should always be a top priority. In this work, we anonymize user information where appropriate to prevent the leakage of identifiable information.
@@ Acknowledgment
We express our gratitude to Myket Corporation for generously providing the anonymous interaction data utilized in this research. We are thankful to Dr. Vahid Rahimian for agreeing to collaborate and for their support throughout the data-sharing process and to Ms. Zahra Eskandari for her diligent efforts in collecting and arranging the data. Additionally, we thank MohammadAmin Fazli for his continuous support and help in facilitating collaboration and data sharing. We thank Weihua Hu, Matthias Fey, Jure Leskovec and Michael Bronstein from the TGB team for their support and discussion. We thank the OGB team for sharing their website template for the TGB website. This research was enabled in part by compute resources, software and technical help provided by Mila. This research was supported by the Canadian Institute for Advanced Research (CIFAR AI chair program), Natural Sciences and Engineering Research Council of Canada (NSERC) Postgraduate Scholarship Doctoral (PGS D) Award and Fonds de recherche du Québec - Nature et Technologies (FRQNT) Doctoral Award.
Appendix A Broader Impact